prove that the function g(x)=ln(x)-1/4 has a solution in the interval (0,5)


1
Expert's answer
2021-07-12T16:45:07-0400

The function "g(x)=\ln (x)-\dfrac14" is considered. The function "g" is continuous in "[1,5]" because it is a sum of continuous functions. Moreover, it changes sign at the extremes of the interval:


"g(1)=\ln (1)-\dfrac14=0-\dfrac14=-\dfrac14<0."

"g(5)=\ln (5)-\dfrac14>0."

Therefore, by Bolzano's theorem, there exists "c\in (1,5)" , such that "g(c)=0" , which proves that "g" has a zero in this interval, and therefore, also in the interval "(0,5)" because "g" is well-defined in this interval.